OpenStudy (moongazer):

how to solve this simple equation? ln[(145-X)/(70-X)] =2*(110-X)/(70-X)

OpenStudy (anonymous):

Don't Keep Me Waitin' Lad!

OpenStudy (moongazer):

there

OpenStudy (moongazer):

what about e^x = x ?

OpenStudy (cj49):

this is the question ryt..(145-X)/(70-X)] =2*(110-X)/(70-X)

OpenStudy (moongazer):

I tried it in wolframalpha but I didn't understand the answer. It should have ln at the beginning

OpenStudy (moongazer):